Find the diameter of a circle whose area as 112 cm² round your answer to the nearest 10th
Answer by Theo(13342)

formula for area of a circle is a = pi * r^2.

a is the area
pi is 3.14.....
r is the radius.
d is the diameter.

the diameter is equal to twice the length of the radius.

a = pi * r^2.
since a = 112, the formula becomes 112 = pi * r^2.
divide both sides of this equation by pi to get:
112/pi = r^2
take the square root of both sides of this equation and solve for r to get:
r = plus or minus sqrt(112/pi)
the radius can't be minus, so you get:
r = sqrt(112/pi)
since diameter = twice the radius, you get:
d = 2 * r which makes d = 2 * sqrt(112)/pi)
solve for d to get d = 11.94164264
round to 1 decimal place to get d = 11.9